Statement of Cash Flows—Indirect Method

The comparative balance sheet of Olson-Jones Industries Inc. for December 31, 20Y2 and 20Y1, is as follows:

Dec. 31, 20Y2 Dec. 31, 20Y1
Assets
Cash $203 $65
Accounts receivable (net) 116 81
Inventories 73 44
Land 166 184
Equipment 93 71
Accumulated depreciation-equipment (25) (13)
Total Assets $626 $432
Liabilities and Stockholders' Equity
Accounts payable (merchandise creditors) $79 $65
Dividends payable 13 -
Common stock, $10 par 41 20
Paid-in capital: Excess of issue price over par—common stock 103 51
Retained earnings 390 296
Total liabilities and stockholders' equity $626 $432

The following additional information is taken from the records:

  1. Land was sold for $45.
  2. Equipment was acquired for cash.
  3. There were no disposals of equipment during the year.
  4. The common stock was issued for cash.
  5. There was a $135 credit to Retained Earnings for net income.
  6. There was a $41 debit to Retained Earnings for cash dividends declared.

a. Prepare a statement of cash flows, using the indirect method of presenting cash flows from operating activities. Use the minus sign to indicate cash out flows, cash payments, decreases in cash, or any negative adjustments.

Olson-Jones Industries, Inc.
Statement of Cash Flows
For the Year Ended December 31, 20Y2
Cash flows from operating activities:
$
Adjustments to reconcile net income to net cash flow from operating activities:
Changes in current operating assets and liabilities:
Net cash flow from operating activities $
Cash flows from investing activities:
$
Net cash flow provided by investing activities
Cash flows from financing activities:
$
Net cash flow provided by financing activities
$
Cash at the beginning of the year
Cash at the end of the year $

b. Was Olson-Jones’s net cash flow from operations more or less than net income?

The source(s) of the difference are:

  1. Gain on the sale of land
  2. Purchase of equipment
  3. Sale of common stock
  4. Changes in current operating assets and liabilities
  5. Depreciation expense
  6. Dividends paid
Answer #1

X fic a. AB $ 135 $ 12 $ 27 $ 35 $ 29 $ 14 Olson-Jones Industries, Inc. Statement of Cash Flows For the Year Ended December 3

Olson-Joness net cash flow from operations was less than net income The source(s) of the difference are: a. Gain on the sale

for formulas and calculations, refer to the image below -

B2 X for a. =25-13 --(45-(184-166)) =81-116 -44-73 =79-65 Olson-Jones Industries, Inc. Statement of Cash Flows For the Year E

In case you have any query, kindly ask in comments.